40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 11 18:34:11 2014, Jenne, Update, LSC, Composite Error Signal for ARms (7) spectra-ungarble.pdfScreenshot-1.pngScanFromOffToOnResonance.pdfScanThroughResonance.pdf
    Reply  Tue Mar 11 19:38:26 2014, manasa, Update, Green Locking, ALS Slow servo settings 
       Reply  Fri Mar 14 10:31:10 2014, jamie, Update, Green Locking, ALS Slow servo settings 
Message ID: 9719     Entry time: Tue Mar 11 18:34:11 2014     Reply to this: 9721
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Composite Error Signal for ARms (7) 

[Nic, Jenne, EricQ, and Koji]

We have used CESAR successfully to bring the Xarm into resonance.  We start with the ALS signal, then as we approach resonance, the error signal is automatically transitioned to 1/sqrt(TRX), and ramped from there to POX, which we use as the PDH signal.

In the first plot, we have several spectra of the CESAR output signal (which is the error signal for the Xarm), at different arm resonance conditions.  Dark blue is the signal when we are locked with the ALS beatnote, far from IR resonance.  Gold is when we are starting to see IR resonance (arm buildup of about 0.03 or more), and we are using the 1/sqrt(TRX) signal for locking.  Cyan is after we have achieved resonance, and are using only the POX PDH signal.  Purple is the same condition as cyan, except that we have also engaged the low frequency boosts (FM 2, 3, 4) in the locking servo.  FM4 is only usable once you are at IR resonance, and locked using the PDH signal.  We see in the plot that our high frequency noise (and total RMS) decreases with each stage of CESAR (ALS, 1/sqrt(TR) and PDH). 

To actually achieve the gold noise level of 1/sqrt(TR), we first had to increase the analog gain by swapping out a resistor on the whitening board. 

 

spectra-ungarble.pdf

The other plots attached are time series data.  For the python plots (last 2), the error signals are calibrated to nanometers, but the dark blue, which is the transmitted power of the cavity, is left in normalized power units (where 1 is full IR resonance). 

In the scan from off resonance to on resonance, around the 58 second mark, we see a glitch when we engage FM4, the strong low frequency boosts.  Around the 75 second mark we turned off any contribution from 1/sqrt(TR), so the noise decreases once we are on pure PDH signal. 

In the scan through the resonance, we see a little more clearly the glitch that happens when we switch from ALS to IR signals, around the 7 and 12 second marks. 

We want to make some changes, so that the transition from ALS to IR signals is more smooth, and not a discrete switch.

 

Attachment 2: Screenshot-1.png  21 kB  Uploaded Tue Mar 11 22:31:40 2014  | Hide | Hide all
Screenshot-1.png
Attachment 3: ScanFromOffToOnResonance.pdf  147 kB  Uploaded Tue Mar 11 23:31:14 2014  | Hide | Hide all
ScanFromOffToOnResonance.pdf
Attachment 4: ScanThroughResonance.pdf  164 kB  Uploaded Tue Mar 11 23:38:10 2014  | Hide | Hide all
ScanThroughResonance.pdf
ELOG V3.1.3-